OpenStudy (anonymous):

The sum of three consecutive integers is 267. What is the largest integer?

OpenStudy (anonymous):

Oh, sorry I should change my words. Consider that if the smallest integer times three, it's product will be 267 if it add what number?

OpenStudy (anonymous):

If it add 3. If the smallest number is x, the second is x+1, the largest is x+2. Add them together, x+x+1+x+2=3x+3 So 267-3=264 264/3=88 88 is the smallest number which is 2 smaller than the largest number. Thus the largest number is 88+2=90 And you can find that 88+89+90=267

OpenStudy (anonymous):

thanks for your help!!!!!

OpenStudy (anonymous):

You are welcome!